# A study of gamma-ray emission from OJ 287 using Fermi-LAT from 2015-2023

### v1

@@ Abstract
We search for gamma-ray emission from OJ287 in the energy range from 0.1-300 GeV during 2015-2023, in coincidence with an extensive observing campaign to monitor the optical flux variability and polarization, as discussed in arXiv:2311.02372. We present results for eight segments in the aforementioned period, with each segment corresponding to an observing season. We find non-zero gamma-ray emission for all the eight segments. The photon and energy flux observed during this period is $\sim 5.2\times 10^{-8}\rm{ph~{}cm^{-2}~{}s^{-1}}$ and $2.75\times 10^{-5}\rm{MeV~{}cm^{-2}~{}s^{-1}}$, respectively, while the SED can be fitted with a power law with a slope of $2.12\pm 0.02$. The observed luminosity is between $10^{45}-10^{46}$ ergs/sec.
@@ 1 Introduction
OJ287 is a blazar located at a redshift of z=0.306 ($D_{L}\sim 1600$ Mpc)  (Sitko and Junkkarinen, 1985). This blazar is known to contain a supermassive black hole binary and hence has been a promising target for gravitational wave astronomy (Valtonen et al., 2021; Gao et al., 2024; Chen and Zhang, 2018; Dey et al., 2018; Desai et al., 2008). Optical observations for this object go back to the 19th century (Sillanpaa et al., 1988). This object has been known to emit quasi-periodic outbursts with a period of approximately 12 years. These bursts are emitted due to the enhanced accretion from the secondary black hole on to the primary (Dey et al., 2019; Valtonen et al., 2021, 2023).
@@
In a recent work, Gupta et al. (2023) reported results of the optical flux variability and polarization measurements using a large network of telescopes (Perkins telescope, 1.5 m “Kanata” telescope, St Petersbug University 70-cm AZT-8 telescope and 40-cm LX-200 telescope, 0.6 m telescope Belogradchik Observatory, Steward observatory) spanning eight observing seasons from September 2015 to October 2023. These observations are part of a long-term multiwavelength campaign of OJ 287 (Gupta et al., 2023). Observations were presented for eight segments for the above duration, where each segment roughly corresponds to one observing cycle. The third segment was further subdivided into two sub-segments labelled as 3A and 3B because of some peculiar variability features in the third segment. Most recently,  Valtonen et al. (2024) also reported a huge optical flare in both $I$-band and $R$-band with TESS and the prompt5 telescope in Chile, on Nov 12, 2021 at 2 am UT, which falls during segment 7. Also, TeV gamma-ray emission has also been detected from OJ287 using Veritas between 1 February and 4 February 2017 (UTC) which falls during Segment 2 (VERITAS Collaboration, 2017), although no details on the duration of the TeV flare are available.
@@
In this work, we search for gamma-ray emission during four flaring periods from 2015-2023 using the Fermi Large Area Telescope (LAT) gamma-ray telescope. The LAT is one of the two instruments onboard this detector. Fermi-LAT is sensitive to high-energy gamma rays from a plethora of astrophysical sources (Ajello et al., 2019). It is a pair-conversion telescope with a field of view of aboud 2.4 sr that is sensitive to photons between the energy range of 30 MeV to around 300 GeV and has been operating continuously since 2008 (Atwood et al., 2009). Fermi-LAT has been monitoring OJ287 since the start of the mission. OJ 287 has been included in the 3FGL (Acero et al., 2015) and 4FGL (Abdollahi et al., 2020) catalogs and has been refered to as J0854.8+2006 in the 4FGL (Ajello et al., 2020). Previous gamma-ray variability studies with Fermi-LAT have been carried out in  Prince et al. (2021a). They found that the isotropic gamma-ray luminosity to be between $10^{47}-10^{48}$ ergs/sec. The average and peak photon flux were found to $\sim 2.65$ and $\sim 10$$\times 10^{-8}\rm{ph.cm^{-2}sec^{-1}}$ in the energy range between 0.1-300 GeV (Prince et al., 2021a). A variation of factor of five was found between the lowest and highest flux states from this analysis. We now present the results from our analysis of Fermi-LAT data between 2015-2023.
@@ 2 Analysis of Fermi-LAT data
We analyzed the Fermi-LAT data using the EasyFermi software (de Menezes, 2022). The Fermi-LAT data from 2015 - 2023 was divided into eight time segments. For each segment, we searched for OJ 287 in the 4FGL-DR3 catalog, within an energy range of 100 MeV to 300 GeV (unless specified) and a search radius of $7.5^{\circ}$. We used 20-time bins for Light curves for Segments 1, 2, 3A, 5, 7, and 8; 8 bins for Segment 3B; 7 bins for Segment 4; and 10 bins for Segment 6 (as larger number of bins have resulted in null data points in some bins). We then proceeded to fit the Spectral Energy Distribution (SED) using a simple power law model.
@@
where $\alpha$ is the spectral index, $N_{0}$ is the normalization and $E_{0}$ is the pivot energy.
@@ 3 Results
We now show the results for the gamma-ray light curves and spectral energy distribution (SED), with and without correction for absorption due to extragalactic background light, and photon flux light curves for OJ 287 during the aforementioned eight segments. A tabular summary of the results is summarized in Table 1, where we show the test statistic (TS) (Mattox et al., 1996; Pasumarti and Desai, 2022; Manna and Desai, 2024), photon flux, energy flux, pivot energy used for normalization, the best-fit spectral parameters and the isotropic gamma-ray luminosity, which has been calculated using Equation 5 of  Prince et al. (2021b).
@@ 3.1 Segment 1 (JD: 2457284 - 2457536)
Figure 1: Segment 1 (MJD : 57283.5 - 57535.5 ) - Light Curve
Figure 2: Segment 1 - SED
Figure 3: Segment 1 - SED Best fit parameters: $\log_{10}N_{0}=-9.51\pm 0.03$ ; $\alpha=2.10\pm 0.04$; $E_{0}=636.6$ MeV
@@ 3.2 Segment 2 (JD: 2457645 - 2457920)
The photon flux light curve, as well as the SED for segment 2, can be found in Fig. 4 and  5, respectively. The average photon and energy flux are given by $(4.07\pm 0.05)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(3.52\pm 0.09)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. Here, we see a factor of four variation between the smallest and largest flux values. We also find a quasi-periodic variation in the light curves with a period of approximately 12-13 days. The posteriors for the best-fit power law parameters can be found in Fig. 6. The SED shows non-zero flux up to 10 GeV and upper limits beyond that.
Figure 4: Segment 2 (MJD : 57644.5 - 57919.5 ) - Light Curve
Figure 5: Segment 2 - SED
Figure 6: Segment 2 - SED Best fit parameters: $\log_{10}N_{0}=-9.94\pm 0.05$ ; $\alpha=1.92\pm 0.05$ ; $E_{0}=720.59$ MeV
@@ 3.3 Segment 3A (JD: 2458016 - 2458142)
Figure 7: Segment 3A (MJD : 58015.5 - 58141.5 ) - Light Curve
Figure 8: Segment 3A - SED
Figure 9: Segment 3A - SED Best fit parameters: $\log_{10}N_{0}=-9.89\pm 0.16$ ; $\alpha=2.19\pm 0.19$ ; $E_{0}=846.77$ MeV
@@ 3.4 Segment 3B (JD: 2458156 - 2458292)
The photon flux light curve and SED for segment 3B can be found in Fig. 10 and  11, respectively. The average photon and energy flux are given by $(3.88\pm 0.67)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(1.86\pm 0.33)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux until 4 GeV and upper limits beyond that, except for one data point near 50 GeV. Similar to segment 3A, we see an approximately constant flux, except for the first data point, which is close to zero. The posteriors for the best-fit spectral parameters can be found in Fig. 12.
Figure 10: Segment 3B (MJD : $5$ 58291.5 ) - Light Curve
Figure 11: Segment 3B - SED
Figure 12: Segment 3B - SED Best fit parameters: $\log_{10}N_{0}=-10.06\pm 0.12$ ; $\alpha=1.97\pm 0.12$ ; $E_{0}=803.72$ MeV
@@ 3.5 Segment 4 (JD: 2458380 - 2458632)
The photon flux light curve and SED for segment 4 can be found in Fig. 13 and  14, respectively. The average photon and energy flux are given by $(4.80\pm 0.78)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.02\pm 0.34)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux until 3 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve shows a signature of quasi-periodic trends, only the period changes from the beginning to the end of the segment. The light curve also shows a factor of five variation between the minimum and maximum flux. The posteriors for the best-fit spectral parameters can be found in Fig. 15.
Figure 13: Segment 4 (MJD : 58379.5 - 58631.5 ) - Light Curve
Figure 14: Segment 4 - SED
Figure 15: Segment 4 - SED Best fit parameters: $\log_{10}N_{0}=-9.90\pm 0.06$ ; $\alpha=2.21\pm 0.10$ ; $E_{0}=545.92$ MeV
@@ 3.6 Segment 5 (JD: 2458755 - 2459009)
Figure 16: Segment 5 (MJD : 58754.5 - 59008.5 ) - Light Curve
Figure 17: Segment 5 - SED
Figure 18: Segment 5 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.05$ ; $\alpha=2.10^{+0.06}_{-0.07}$ ; $E_{0}=921.45$ MeV
@@ 3.7 Segment 6 (JD: 2459108 - 2459338)
The photon flux light curve and SED for segment 6 can be found in Fig. 19 and  20, respectively. The average photon and energy flux are given by $(3.72\pm 0.67)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.06\pm 0.33)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 10 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve mainly shows a constant flux, except for the first and third point which depict very small values and then rises towards the end of the segment. The posteriors for the best-fit spectral parameters can be found in Fig. 21.
Figure 19: Segment 6 (MJD : 59107.5 - 59337.5 ) - Light Curve
Figure 20: Segment 6 - SED
Figure 21: Segment 6 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.06$ ; $\alpha=2.04\pm 0.08$ ; $E_{0}=745.69$ MeV
@@ 3.8 Segment 7 (JD: 2459520 - 2459750)
The photon flux light curve and SED for segment 7 can be found in Fig. 22 and  23, respectively. The average photon and energy flux are given by $(2.06\pm 0.25)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(1.95\pm 0.27)\times 10^{-5}$ $\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 20 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve shows a sinusoidal trend near the start of the segment until MJD of 59600 followed by zero flux for the next 50 days. The light curve then shows a constant value until the end of the segment. We also do not find any correlation between the optical flare in $I$ and $R$ band reported in  Valtonen et al. (2024) and the gamma-ray light curve. The posteriors for the best-fit spectral parameters can be found in Fig. 24.
Figure 22: Segment 7 (MJD : 59519.5 - 59749.5 ) - Light Curve. The dashed vertical line corresponds to optical flare detected in $I$ and $R$ band as reported in  Valtonen et al. (2024).
Figure 23: Segment 7 - SED
Figure 24: Segment 7 - SED Best fit parameters: $\log_{10}N_{0}=-10.20^{+0.13}_{-0.12}$ ; $\alpha=1.85^{+0.12}_{-0.11}$ ; $E_{0}=1081.02$ MeV

### v2

@@ Abstract
We search for gamma-ray emission from OJ287 in the energy range from 0.1-300 GeV during 2015-2023, in coincidence with an extensive observing campaign to monitor the optical flux variability and polarization, as discussed in arXiv:2311.02372. We present results for eight segments in the aforementioned period, with each segment corresponding to an observing season. We report non-zero gamma-ray flux (with $>10\sigma$ significance) for all the eight segments. The photon and energy flux observed during this period is $\sim 5.2\times 10^{-8}\rm{ph~{}cm^{-2}~{}s^{-1}}$ and $2.75\times 10^{-5}\rm{MeV~{}cm^{-2}~{}s^{-1}}$, respectively, while the SED can be fitted with a power law with a slope of $2.12\pm 0.02$. The observed luminosity is between $10^{45}-10^{46}$ ergs/sec. Similar to previous works we find a hard cutoff in the gamma-ray spectrum at around 20 GeV. We do not observe any correlations between the peaks in the gamma ray light curves and corresponding optical light curves in $I$ and $R$ bands, as well as the X-ray light curves from SWIFT between 0.3-10 keV.
@@ 1 Introduction
OJ287 is a BL object located at a redshift of z=0.306 ($D_{L}\sim 1600$ Mpc)  (Sitko and Junkkarinen, 1985). This object is known to contain a supermassive black hole binary and hence has been a promising target for studies in gravitational wave astronomy (Valtonen et al., 2021; Gao et al., 2024; Chen and Zhang, 2018; Dey et al., 2018; Desai et al., 2008). Optical observations for this object go back to the 19th century (Sillanpaa et al., 1988; Hudec et al., 2013). This object has been known to emit quasi-periodic outbursts with a period of approximately 12 years. These optical bursts are emitted due to the enhanced accretion from the secondary black hole onto the primary (Dey et al., 2019; Valtonen et al., 2021, 2023).
@@
In a recent work, Gupta et al. (2023) reported results related to optical flux variability and polarization measurements using a large network of telescopes (Perkins telescope, 1.5 m “Kanata” telescope, St Petersburg University 70-cm AZT-8 telescope and 40-cm LX-200 telescope, 0.6 m telescope at the Belogradchik Observatory, Steward Observatory telescope) spanning eight observing seasons from September 2015 to October 2023. These observations are part of a long-term multiwavelength campaign of OJ 287 (Gupta et al., 2023). Observations were presented for eight segments for the above duration, where each segment roughly corresponds to one observing cycle. The third segment was further subdivided into two sub-segments labelled as 3A and 3B because of some peculiar variability features in the third segment. Most recently,  Valtonen et al. (2024) also reported a huge optical flare in November 2021 in both $I$-band and $R$-band with TESS and the prompt5 telescope in Chile as part of the Krakow quasar monitoring campaign, which falls during segment 7. Also, TeV gamma-ray emission has also been detected from OJ287 using VERITAS between 1 February and 4 February 2017 (UTC), which falls during Segment 2 (VERITAS Collaboration, 2017), although no details on the duration of the TeV flare are available. Finally this object has also been monitored in X-rays part of the MOMO project (Komossa et al., 2023a).
@@
In this work, we search for gamma-ray emission during the eight aforementioned segments from 2015-2023 using the Fermi Large Area Telescope (LAT) gamma-ray telescope. The LAT is one of the two instruments onboard this detector. Fermi-LAT is sensitive to high-energy gamma rays from a plethora of astrophysical sources (Ajello et al., 2019). It is a pair-conversion telescope with a field of view of about 2.4 sr that is sensitive to photons between the energy range of 30 MeV to around 300 GeV and has been operating continuously since 2008 (Atwood et al., 2009). Fermi-LAT has been monitoring OJ287 since the start of the mission. OJ 287 has been included in the 3FGL (Acero et al., 2015) and 4FGL (Abdollahi et al., 2020) catalogs and has been referred to as J0854.8+2006 in the 4FGL (Ajello et al., 2020). Previous gamma-ray variability studies with Fermi-LAT have been reported in a number of works (Abdo et al., 2009; Neronov and Vovk, 2011; Kushwaha et al., 2013; Prince et al., 2021a; Komossa et al., 2022).  Prince et al. (2021a) found that the isotropic gamma-ray luminosity to be between $10^{47}-10^{48}$ ergs/sec. The average and peak photon flux were found to be $\sim 2.65$ and $\sim 10$$\times 10^{-8}\rm{ph.cm^{-2}sec^{-1}}$ in the energy range between 0.1-300 GeV (Prince et al., 2021a). This analysis found a variation of a factor of five between the lowest and highest flux states. We now present the results from our analysis of Fermi-LAT data between 2015-2023.
This manuscript is structured as follows. Our analysis of the Fermi-LAT data is discussed in Sect 2. Comparison with other multi-wavelength flares for OJ 287 is discussed in Sect. 4.2. We conclude in Sect. 5.
@@ 2 Analysis of Fermi-LAT data
We analyzed the Fermi-LAT data using the EasyFermi software (de Menezes, 2022), which is based on both FermiTools and Fermipy packages (Wood et al., 2017). The Fermi-LAT data from 2015 - 2023 was divided into eight time segments. For each segment, we searched for OJ 287 in the 4FGL-DR3 catalog, within an energy range of 100 MeV to 300 GeV (unless specified) and a search radius of $7.5^{\circ}$. We used 20-time bins for light curves for Segments 1, 2, 3A, 5, 7, and 8; 8 bins for Segment 3B; 7 bins for Segment 4; and 10 bins for Segment 6 (as larger number of bins have resulted in null data points in some bins). We then proceeded to fit the Spectral Energy Distribution (SED) using a simple power law model and also a log-parabola model. The power-law spectrum is defined as follows:
@@
where $\alpha$ is the spectral index, $N_{0}$ is the normalization and $E_{0}$ is the pivot energy. Similarly, the log-parabola spectrum is defined according to:
$dN/dE=N_{0}\left(\frac{E}{E_{0}}\right)^{-\alpha-\beta\log{\left(\frac{E}{E_{0}}\right)}}$ (2)
@@ 3 Results
We now show the results for the gamma-ray light curves and SED, with and without correction for absorption due to extragalactic background light, and photon flux light curves for OJ 287 during the aforementioned eight segments. A tabular summary of the results is summarized in Table 1, where we show the test statistic (TS) (Mattox et al., 1996; Pasumarti and Desai, 2022; Manna and Desai, 2024), photon flux, energy flux, and the isotropic gamma-ray luminosity, which has been calculated using Equation 5 of  Prince et al. (2021b). The best-fit spectral parameters for both the power-law and log-parabola model can be found in Table 2. We also confirmed that that the value of “Fit Quality” reported by EasyFermi is equal to three for all the analyzed segments, indicating the fit is excellent and there is no problem in the convergence of the log likelihood de Menezes (2022).
@@ 3.1 Segment 1 (JD: 2457284 - 2457536)
Figure 1: Segment 1 (MJD : 57283.5 - 57535.5 ) - Light Curve; binwidth = 12 days.
Figure 2: Segment 1 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 3: Segment 1 - SED Best fit parameters: $\log_{10}N_{0}=-9.51\pm 0.03$ ; $\alpha=2.10\pm 0.04$; $E_{0}=636.6$ MeV.
@@ 3.2 Segment 2 (JD: 2457645 - 2457920)
The photon flux light curve, as well as the SED for segment 2, can be found in Fig. 4 and  5, respectively. The average photon and energy flux are given by $(4.07\pm 0.05)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(3.52\pm 0.09)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. Here, we see a factor of four variation between the smallest and largest flux values. The posteriors for the best-fit power law parameters can be found in Fig. 6. The SED shows non-zero flux up to 10 GeV and upper limits beyond that.
Figure 4: Segment 2 (MJD : 57644.5 - 57919.5 ) - Light Curve; binwidth = 13.75 days
Figure 5: Segment 2 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 6: Segment 2 - SED Best fit parameters: $\log_{10}N_{0}=-9.94\pm 0.05$ ; $\alpha=1.92\pm 0.05$ ; $E_{0}=720.59$ MeV.
@@ 3.3 Segment 3A (JD: 2458016 - 2458142)
Figure 7: Segment 3A (MJD : 58015.5 - 58141.5 ) - Light Curve; binwidth = 6.3 days
Figure 8: Segment 3A - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 9: Segment 3A - SED Best fit parameters: $\log_{10}N_{0}=-9.89\pm 0.16$ ; $\alpha=2.19\pm 0.19$ ; $E_{0}=846.77$ MeV.
@@ 3.4 Segment 3B (JD: 2458156 - 2458292)
The photon flux light curve and SED for segment 3B can be found in Fig. 10 and  11, respectively. The average photon and energy flux are given by $(3.88\pm 0.67)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(1.86\pm 0.33)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux until 4 GeV and upper limits beyond that, except for one data point near 50 GeV. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The posteriors for the best-fit spectral parameters can be found in Fig. 12.
Figure 10: Segment 3B (MJD : $5$ 58291.5 ) - Light Curve ; binwidth = 17 days
Figure 11: Segment 3B - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 12: Segment 3B - SED Best fit parameters: $\log_{10}N_{0}=-10.06\pm 0.12$ ; $\alpha=1.97\pm 0.12$ ; $E_{0}=803.72$ MeV.
@@ 3.5 Segment 4 (JD: 2458380 - 2458632)
The photon flux light curve and SED for segment 4 can be found in Fig. 13 and  14, respectively. The average photon and energy flux are given by $(4.80\pm 0.78)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.02\pm 0.34)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux until 3 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve also shows a factor of five variation between the minimum and maximum flux. The posteriors for the best-fit spectral parameters can be found in Fig. 15.
Figure 13: Segment 4 (MJD : 58379.5 - 58631.5 ) - Light Curve ; binwidth = 36 days
Figure 14: Segment 4 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 15: Segment 4 - SED Best fit parameters: $\log_{10}N_{0}=-9.90\pm 0.06$ ; $\alpha=2.21\pm 0.10$ ; $E_{0}=545.92$ MeV.
@@ 3.6 Segment 5 (JD: 2458755 - 2459009)
Figure 16: Segment 5 (MJD : 58754.5 - 59008.5 ) - Light Curve ; binwidth = 12.7 days
Figure 17: Segment 5 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 18: Segment 5 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.05$ ; $\alpha=2.10^{+0.06}_{-0.07}$ ; $E_{0}=921.45$ MeV.
@@ 3.7 Segment 6 (JD: 2459108 - 2459338)
The photon flux light curve and SED for segment 6 can be found in Fig. 19 and  20, respectively. The average photon and energy flux are given by $(3.72\pm 0.67)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.06\pm 0.33)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively.
From the SED, we find non-zero flux up to approximately 10 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve mainly shows a constant flux, except for the first and third points, which depict very small values and then rise towards the end of the segment. The posteriors for the best-fit spectral parameters can be found in Fig. 21.
Figure 19: Segment 6 (MJD : 59107.5 - 59337.5 ) - Light Curve ; binwidth = 23 days
Figure 20: Segment 6 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 21: Segment 6 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.06$ ; $\alpha=2.04\pm 0.08$ ; $E_{0}=745.69$ MeV.
@@ 3.8 Segment 7 (JD: 2459520 - 2459750)
The photon flux light curve and SED for segment 7 can be found in Fig. 22 and  23, respectively. The average photon and energy flux are given by $(2.06\pm 0.25)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(1.95\pm 0.27)\times 10^{-5}$ $\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 20 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve shows a sinusoidal trend near the start of the segment until MJD of 59600 followed by upper limits for the next 50 days. The light curve then shows a constant value until the end of the segment. We also do not find any correlation between the optical flare in $I$ and $R$ band reported in  Valtonen et al. (2024) and the gamma-ray light curve. The posteriors for the best-fit spectral parameters can be found in Fig. 24.
Figure 22: Segment 7 (MJD : 59519.5 - 59749.5 ) - Light Curve ; binwidth = 11.5 days. The dashed vertical line corresponds to the start of the optical flare detected in $I$ and $R$ band as reported in  Valtonen et al. (2024).
Figure 23: Segment 7 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 24: Segment 7 - SED Best fit parameters: $\log_{10}N_{0}=-10.20^{+0.13}_{-0.12}$ ; $\alpha=1.85^{+0.12}_{-0.11}$ ; $E_{0}=1081.02$ MeV.